# Chapter 0 Strong CP problem, theta term and QCD topological properties

### v1

@@ 1 Introduction
While this topological $\theta$ term would be completely irrelevant at the classical level (being it the integral of a four-divergence), it has actually physical consequences in the quantum theory. Indeed the path-integral which is expected to non-perturbatively define QCD as a Quantum Field Theory can be decomposed into a sum over all the topological sectors (i. e. the subsets of the gauge field space with fixed values of the topological charge) and the introduction of a $\theta$ term changes the relative weight of the different sectors, thus introducing a non-trivial dependence on this parameter in all physical observables.
@@ 1 Why $\theta$-dependence
The $\theta$-dependence of four dimensional gauge theories is typically introduced using the semiclassical approximation and instantons, see, e. g., ColemanBook ; ShifmanBook ; Weinberg2Book ; ZinnJustinBook . A canonical approach can however also be used, which emphasizes the role of locality and of the topology of the gauge group (as opposed to the topology of field configurations with finite Euclidean action), clarifying that $\theta$-dependence is a physical property of the theory independent of the semiclassical approximation. Such an approach has been developed in Jackiw:1979ur ; Jackiw85 , and its influence can be seen, e. g., in the presentations of WeinbergEBook and, particularly, of SmilgaBook .
Before addressing the case of four-dimensional gauge theories, let us start by discussing in some detail what is arguably the technically simplest system characterized by $\theta$-dependence, i. e., a quantum particle moving on a circumference. This system, despite its extreme simplicity, displays several key features which have analogues in four-dimensional gauge theories Jackiw:1979ur ; SmilgaBook ; SchulmanBook ; Bietenholz:1997kr ; Gaiotto:2017yup ; Bonati:2017woi . If we denote by $\phi$ the angle on the circumference, the Hamiltonian of a quantum particle of mass $m$ moving on a circumference of radius $R$ is
@@
$_{\theta^{\prime}}\langle\phi_{f},t_{f}|\phi_{i},t_{i}\rangle_{\theta}=2\pi\delta(\theta-\theta^{\prime})\sum_{n}\psi_{\theta,n}(\phi_{f})\psi_{\theta,n}^{*}(\phi_{i})\mathrm{e}^{-\mathrm{i}E^{(\theta)}_{n}(t_{f}-t_{i})/\hbar}=2\pi\delta(\theta-\theta^{\prime})\left.\langle\phi_{f},t_{f}|\phi_{i},t_{i}\rangle\right|_{\theta}\ .$ (7)
@@
$\left.\langle\phi_{f},t_{f}|\phi_{i},t_{i}\rangle\right|_{\theta}$ $=\int_{0}^{2\pi}\mathrm{d}\theta^{\prime}2\pi\delta(\theta-\theta^{\prime})\sum_{n}\psi_{\theta^{\prime},n}(\phi_{f})\psi_{\theta^{\prime},n}^{*}(\phi_{i})\mathrm{e}^{-\mathrm{i}E^{(\theta^{\prime})}_{n}(t_{f}-t_{i})/\hbar}=\int_{0}^{2\pi}\mathrm{d}\theta^{\prime}\sum_{Q=-\infty}^{\infty}\mathrm{e}^{-\mathrm{i}Q(\theta-\theta^{\prime})}\sum_{n}\psi_{\theta^{\prime},n}(\phi_{f})\psi_{\theta^{\prime},n}^{*}(\phi_{i})\mathrm{e}^{-\mathrm{i}E^{(\theta^{\prime})}_{n}(t_{f}-t_{i})/\hbar}=$ (9)
$=\sum_{Q=-\infty}^{\infty}\mathrm{e}^{-\mathrm{i}Q\theta}\int_{0}^{2\pi}\mathrm{d}\theta^{\prime}\sum_{n}\psi_{\theta^{\prime},n}(\phi_{f}+2\pi Q)\psi_{\theta^{\prime},n}^{*}(\phi_{i})\mathrm{e}^{-\mathrm{i}E^{(\theta^{\prime})}_{n}(t_{f}-t_{i})/\hbar}=\sum_{Q=-\infty}^{\infty}\mathrm{e}^{-\mathrm{i}Q\theta}\langle\phi_{f}+2\pi Q,t_{f}|\phi_{i},t_{i}\rangle=$
$=\sum_{Q=-\infty}^{\infty}\mathrm{e}^{-\mathrm{i}Q\theta}\int_{\phi(t_{i})=\phi_{i}}^{\phi(t_{f})=\phi_{f}+2\pi Q}\mathcal{D}\phi(t)\,\exp\left(\,\frac{\mathrm{i}}{\hbar}\int_{t_{i}}^{\,t_{f}}L(\phi,\dot{\phi})\mathrm{d}t\right)=\int_{\phi(t_{i})\,=\,\phi_{i}}^{\phi(t_{f})\,=\,\phi_{f}}\mathcal{D}\phi(t)\,\exp\left(\,\frac{\mathrm{i}}{\hbar}\int_{t_{i}}^{\,t_{f}}\left(L(\phi,\dot{\phi})-\hbar\frac{\theta}{2\pi}\dot{\phi}\right)\mathrm{d}t\right)\ ,$
and we thus obtain the Lagrangian
@@
The last term of the Lagrangian $L_{\theta}$ has the typical form of a $\theta$ term: it is a term of quantum origin (as signaled by the explicit presence of $\hbar$ in the Lagrangian) which is a total derivative with respect to time, hence it has no effect on the classical equation of motion of the system, still it has non trivial effects in the quantum theory. To verify the last statement, i. e. the effect of the $\theta$ term in the quantum theory, there are (at least) two ways. The first, more direct, one consists in explicitly computing the energy spectrum of the model in the free case $V=0$: it is indeed trivial to verify that in this case the energy spectrum explicitly depends on $\theta$, and it is given by
@@
The second method uses the path-integral formulation: the $\theta$ term contributes to the propagator from the state $|\phi_{i}\rangle$ at time $t_{i}$ to the state $|\phi_{f}\rangle$ at time $t_{f}$ with
@@
The effect of the $\theta$ term is thus of topological origin, as can be seen by noting that, for fixed $\phi_{i}(t_{i})$ and $\phi_{f}(t_{f})$, the contributions of two different paths have a relative weight $\exp(-\mathrm{i}\theta\Delta Q)$, where $\Delta Q$ is the difference between the winding numbers along the circumference of the two paths. For the particular case of the quantum particle moving on a circumference we can also note that $\theta$ enters the Lagrangian $L_{\theta}$ in the same way as a constant vector potential directed along the $\phi$ direction:
@@
We are now ready to discuss the case of four dimensional gauge theories, considering for the sake of the simplicity the case of the gauge group SU($N$) and using natural units. The Lagrangian density of the pure glue theory is
@@
All gauge transformations that will be used in the following will be assumed to be local, i. e. $\Omega(\mathbf{x})\to 1$ sufficiently fast as $|\mathbf{x}|\to\infty$ (we denote by $\mathbf{x}$ the spatial components of $x$, at fixed time). Infinitesimal gauge transformations, in particular, have to satisfy $\chi^{a}(\mathbf{x})\to 0$ as $|\textbf{x}|\to\infty$. If we now denote by $\Psi[\mathbf{A}]$ the wavefunction of the system (with A denoting collectively the spatial components $A_{j}(\mathbf{x})$ for a fixed time), the variation of $\Psi[\mathbf{A}]$ due to an infinitesimal gauge transformation is (integrating by part and using $P^{a\,j}(x)=-\mathrm{i}\frac{\delta}{\delta A_{j}^{a}(x)}$)
@@
Since $\chi^{a}(\mathbf{x})$ approaches zero for large x, the previous integral is well defined, and we see that the constraint in Eq. (16) implements gauge invariance under infinitesimal local gauge transformations. If matter is coupled to the gauge field, a further term appears in Eq. (16), related to the variation with respect to $A^{0}$ of the matter part of the Lagrangian density. The resulting constraint once again implements gauge invariance, this time under infinitesimal local gauge transformations of both gauge and matter fields.
@@
We are now almost ready to adapt the reasoning carried out in Eq. (9) to the case of four dimensional gauge theories. The last ingredient needed is a relation between the winding number $N(\Omega)$ and four dimensional integrals. To this purpose we introduce the (non gauge invariant) Chern-Simons current
$j_{CS}^{\mu}=\frac{g^{2}}{16\pi^{2}}\epsilon^{\,\mu\alpha\beta\gamma}\mathrm{Tr}\left(F_{\alpha\beta}A_{\gamma}+\frac{2\mathrm{i}g}{3}A_{\alpha}A_{\beta}A_{\gamma}\right)\ ,$ (23)
@@
$q\equiv\partial_{\mu}j_{CS}^{\mu}=\frac{g^{2}}{16\pi^{2}}\mathrm{Tr}\left(F_{\mu\nu}\tilde{F}^{\mu\nu}\right)=\frac{g^{2}}{32\pi^{2}}\epsilon^{\,\mu\nu\alpha\beta}F^{a}_{\mu\nu}F^{a}_{\alpha\beta}\ ,$ (24)
@@
$W(\mathbf{A})=\int j_{CS}^{0}(t,\textbf{x})\mathrm{d}^{3}\textbf{x}\ .$ (25)
Using the four dimensional divergence theorem and assuming the spatial components of the Chern-Simons current to be trivial at spatial infinity (e. g. due to some fixed boundary conditions) we have
@@
$\left.\langle\mathbf{A}_{f},t_{f}|\textbf{A}_{i},t_{i}\rangle\right|_{\theta}$ $=\int_{0}^{2\pi}\mathrm{d}\theta^{\prime}\sum_{\nu=-\infty}^{\infty}\mathrm{e}^{-\mathrm{i}\nu(\theta-\theta^{\prime})}\sum_{n}\Psi_{\theta^{\prime},n}(\mathbf{A}_{f})\Psi_{\theta^{\prime},n}(\textbf{A}_{i})\mathrm{e}^{-\mathrm{i}E_{n}^{(\theta^{\prime})}(t_{f}-f_{i})}=$ (28)
$=\sum_{\nu=-\infty}^{\infty}\mathrm{e}^{-\mathrm{i}\nu\theta}\int_{0}^{2\pi}\mathrm{d}\theta^{\prime}\sum_{n}\Psi_{\theta^{\prime},n}(\mathbf{A}_{f}^{\Omega_{1}^{\nu}})\Psi_{\theta^{\prime},n}(\textbf{A}_{i})\mathrm{e}^{-\mathrm{i}E_{n}^{(\theta^{\prime})}(t_{f}-f_{i})}=\sum_{\nu=-\infty}^{\infty}\mathrm{e}^{-\mathrm{i}\nu\theta}\langle\mathbf{A}_{f}^{\Omega_{1}^{\nu}},t_{f}|\textbf{A}_{i},t_{i}\rangle=$
$=\sum_{\nu=-\infty}^{\infty}\mathrm{e}^{-\mathrm{i}\nu\theta}\int_{\mathbf{A}(t_{i})=\textbf{A}_{i}}^{\textbf{A}(t_{f})=\textbf{A}_{f}^{\Omega_{1}^{\nu}}}\mathcal{D}A(t,x)\,\exp\left(\mathrm{i}\int_{t_{i}}^{t_{f}}\mathrm{d}t\int\mathrm{d}x\mathcal{L}\right)\propto\int_{\mathbf{A}(t_{i})=\textbf{A}_{i}}^{\textbf{A}(t_{f})=\textbf{A}_{f}}\mathcal{D}A(t,x)\,\exp\left(\mathrm{i}\int_{t_{i}}^{t_{f}}\mathrm{d}t\int\mathrm{d}x\left[\mathcal{L}+\theta q\right]\right)\ ,$
@@ 2 The strong CP problem
The Lagrangian density of QCD, taking into account also the matter fields and the $\theta$ term, is thus
@@
where $D_{\mu}=\partial_{\mu}-igA_{\mu}^{a}T^{a}$ is the covariant derivative in the fundamental representation of the gauge group and “f” is a flavor index. It is well known that the first two terms of this Lagrangian density (just like in Quantum Electrodynamics) are invariant under both parity (P) and charge conjugation (C), however the $\theta$ term explicitly breaks P and CP for generic values of the $\theta$ angle, and this is essentially due to the presence of the Levi-Civita totally antisymmetric tensor. If we introduce the three dimensional chromoelectric and chromomagnetic fields in analogy with classical electromagnetism we have indeed that $q\propto\mathbf{E}^{a}\cdot\mathbf{B}^{a}$, and since the chromomagnetic field is even under parity transformation we see that the $\theta$ term is odd under both P and CP. Invariance under parity and CP is clearly recovered for $\theta=0$, but also for $\theta=\pi$: since $\theta$ is defined up to $2\pi$, we have indeed $\pi\sim-\pi$. Note that at $\theta=\pi$ the discrete symmetries are however expected to be spontaneously broken, since at $\theta=\pi$ a level crossing happens, just like in Eq. (11) (see also the following discussion on $\theta$-dependence in chiral perturbation theory in Sec. 2 and Witten:1980sp ; Gaiotto:2017yup ).
@@
To discuss why this is not possible when considering the whole Standard Model we have to introduce one last ingredient: the chiral anomaly Adler:1969gk ; Bell:1969ts . If we consider a four dimensional gauge theory coupled to a single flavor of quark, the chiral rotation
@@
This relation implies that we can reabsorb $\theta$ in the mass term at the expense of using the complex mass $m\mathrm{e}^{-i\theta}$. Note that this does not change the invariance properties of the theory, since a complex mass is not invariant under P and CP transformation, but gives us the possibility of studying $\theta$-dependence using chiral perturbation theory if light quarks are present (and this is the rationale behind Eq. (31), since $m_{\pi}^{2}\propto m$). An immediate consequence of this fact is that there is no $\theta$-dependence at all if $m=0$, since in this case we can shift $\theta$ at will without affecting the mass term of the Lagrangian. When both the $\theta$ term and a complex mass are present, we can define the physical $\theta$ angle as the combination $\theta_{phys}=\theta-\arg m$, which is invariant under chiral transformations and corresponds to the value of the $\theta$ angle when the fermion mass is real and positive. The previous discussion can be generalized to the case of a generic number of flavors with mass term
@@
where $M$ is a $N_{f}\times N_{f}$ matrix and the physical $\theta$ angle is in this case given by
$\theta_{phys}=\theta-\arg\det M\ .$ (37)
The previous statement regarding the upper limit of the $\theta$ angle should thus more precisely be written as $|\theta_{phys}|\lesssim 10^{-10}$, and the presence of $\theta_{phys}$ shows why we can not simply fix to zero the $\theta$ angle in the Standard Model: the mass matrix originates from the electroweak sector, and since in the electroweak sector P and CP are not conserved, there is no way to fix $\theta_{phys}=0$ using symmetry arguments. The fact that $|\theta_{phys}|$ is so small or, equivalently, the absence of P and CP violations in the strong sector of the Standard Model, constitutes the so called strong CP problem. As noted before there is no $\theta$-dependence at all if at least one quark is massless ($\arg\det M$ is not defined if $M$ is singular). In the past, a massless up quarks has thus been suggested as a simple solution of the strong CP problem, however this is today known to be inconsistent with experimental and lattice QCD results FlavourLatticeAveragingGroupFLAG:2021npn ; Sanfilippo:2015era ; ParticleDataGroup:2024cfk (see also Alexandrou:2020bkd ).

### v2

@@ 1 Introduction
While this topological $\theta$-term would be completely irrelevant at the classical level (being it the integral of a four-divergence), it has actually physical consequences in the quantum theory. Indeed the path-integral which is expected to non-perturbatively define QCD as a Quantum Field Theory can be decomposed into a sum over all the topological sectors (i. e. the subsets of the gauge field space with fixed values of the topological charge) and the introduction of a $\theta$-term changes the relative weight of the different sectors, thus introducing a non-trivial dependence on this parameter in all physical observables.
@@ 1 Why $\theta$-dependence
The $\theta$-dependence of four-dimensional gauge theories is typically introduced using the semiclassical approximation and instantons, see, e. g., ColemanBook ; ShifmanBook ; Weinberg2Book ; ZinnJustinBook . A canonical approach can however also be used, which emphasizes the role of locality and of the topology of the gauge group (as opposed to the topology of field configurations with finite Euclidean action), clarifying that $\theta$-dependence is a physical property of the theory independent of the semiclassical approximation. Such an approach has been developed in Jackiw:1979ur ; Jackiw85 , and its influence can be seen, e. g., in the presentations of WeinbergEBook and, particularly, of SmilgaBook .
Before addressing the case of four-dimensional gauge theories, let us start by discussing in some detail what is arguably the technically simplest system characterized by $\theta$-dependence, i. e., a quantum particle moving on a circumference. This system, despite its extreme simplicity, displays several key features which have analogues in four-dimensional gauge theories Jackiw:1979ur ; Rothe:1977bv ; SmilgaBook ; SchulmanBook ; Bietenholz:1997kr ; Gaiotto:2017yup ; Bonati:2017woi . If we denote by $\phi$ the angle on the circumference, the Hamiltonian of a quantum particle of mass $m$ moving on a circumference of radius $R$ is
@@
$_{\theta^{\prime}}\langle\phi_{f},t_{f}|\phi_{i},t_{i}\rangle_{\theta}=\delta(\theta-\theta^{\prime})\sum_{n}\psi_{\theta,n}(\phi_{f})\psi_{\theta,n}^{*}(\phi_{i})\mathrm{e}^{-\mathrm{i}E^{(\theta)}_{n}(t_{f}-t_{i})/\hbar}=\delta(\theta-\theta^{\prime})\left.\langle\phi_{f},t_{f}|\phi_{i},t_{i}\rangle\right|_{\theta}\ .$ (7)
@@
$\left.\langle\phi_{f},t_{f}|\phi_{i},t_{i}\rangle\right|_{\theta}=\int_{0}^{2\pi}\mathrm{d}\theta^{\prime}\delta(\theta-\theta^{\prime})\sum_{n}\psi_{\theta^{\prime},n}(\phi_{f})\psi_{\theta^{\prime},n}^{*}(\phi_{i})\mathrm{e}^{-\mathrm{i}E^{(\theta^{\prime})}_{n}(t_{f}-t_{i})/\hbar}=\frac{1}{2\pi}\int_{0}^{2\pi}\mathrm{d}\theta^{\prime}\sum_{Q=-\infty}^{\infty}\mathrm{e}^{-\mathrm{i}Q(\theta-\theta^{\prime})}\sum_{n}\psi_{\theta^{\prime},n}(\phi_{f})\psi_{\theta^{\prime},n}^{*}(\phi_{i})\mathrm{e}^{-\mathrm{i}E^{(\theta^{\prime})}_{n}(t_{f}-t_{i})/\hbar}=$ (9)
$\qquad=\frac{1}{2\pi}\sum_{Q=-\infty}^{\infty}\mathrm{e}^{-\mathrm{i}Q\theta}\int_{0}^{2\pi}\mathrm{d}\theta^{\prime}\sum_{n}\psi_{\theta^{\prime},n}(\phi_{f}+2\pi Q)\psi_{\theta^{\prime},n}^{*}(\phi_{i})\mathrm{e}^{-\mathrm{i}E^{(\theta^{\prime})}_{n}(t_{f}-t_{i})/\hbar}=\frac{1}{2\pi}\sum_{Q=-\infty}^{\infty}\mathrm{e}^{-\mathrm{i}Q\theta}\langle\phi_{f}+2\pi Q,t_{f}|\phi_{i},t_{i}\rangle=$
$\qquad=\frac{1}{2\pi}\sum_{Q=-\infty}^{\infty}\mathrm{e}^{-\mathrm{i}Q\theta}\int_{\phi(t_{i})=\phi_{i}}^{\phi(t_{f})=\phi_{f}+2\pi Q}\mathcal{D}\phi(t)\,\exp\left(\,\frac{\mathrm{i}}{\hbar}\int_{t_{i}}^{\,t_{f}}L(\phi,\dot{\phi})\mathrm{d}t\right)\propto\int_{\phi(t_{i})\,=\,\phi_{i}}^{\phi(t_{f})\,=\,\phi_{f}}\mathcal{D}\phi(t)\,\exp\left(\,\frac{\mathrm{i}}{\hbar}\int_{t_{i}}^{\,t_{f}}\left(L(\phi,\dot{\phi})-\hbar\frac{\theta}{2\pi}\dot{\phi}\right)\mathrm{d}t\right)\ ,$
where in the last step $\phi_{i}$ and $\phi_{f}$ are defined modulo $2\pi$, the path-integral is performed by summing on all path topologies and we used $\exp(-\mathrm{i}\frac{\theta}{2\pi}\int_{t_{i}}^{t_{f}}\dot{\phi}\mathrm{d}t)=\mathrm{e}^{-\mathrm{i}Q\theta}\mathrm{e}^{-\mathrm{i}\frac{\theta}{2\pi}(\phi_{f}-\phi_{i})}$. We thus obtain the Lagrangian
@@
The last term of the Lagrangian $L_{\theta}$ has the typical form of a $\theta$ term: it is a term of quantum origin (as signaled by the explicit presence of $\hbar$ in the Lagrangian) which is a total derivative with respect to time, hence it has no effect on the classical equation of motion of the system, still it has non trivial effects in the quantum theory. To verify the last statement, i. e. the effect of the $\theta$-term in the quantum theory, there are (at least) two ways. The first, more direct, one consists in explicitly computing the energy spectrum of the model in the free case $V=0$: it is indeed trivial to verify that in this case the energy spectrum explicitly depends on $\theta$, and it is given by
@@
The second method uses the path-integral formulation: the $\theta$-term contributes to the propagator from the state $|\phi_{i}\rangle$ at time $t_{i}$ to the state $|\phi_{f}\rangle$ at time $t_{f}$ with
@@
The effect of the $\theta$-term is thus of topological origin, as can be seen by noting that, for fixed $\phi_{i}(t_{i})$ and $\phi_{f}(t_{f})$, the contributions of two different paths have a relative weight $\exp(-\mathrm{i}\theta\Delta Q)$, where $\Delta Q$ is the difference between the winding numbers along the circumference of the two paths. Note that the winding number of two open (i.e., with $\phi_{f}\neq\phi_{i}$) paths is not well defined, but their difference is. For the particular case of the quantum particle moving on a circumference we can also note that $\theta$ enters the Lagrangian $L_{\theta}$ in the same way as a constant vector potential directed along the $\phi$ direction:
@@
We are now ready to discuss the case of four-dimensional gauge theories, considering for the sake of the simplicity the case of the gauge group SU($N$) and using natural units. The Lagrangian density of the pure glue theory is
@@
All gauge transformations that will be used in the following will be assumed to be local, i. e. $\Omega(\mathbf{x})\to 1$ sufficiently fast as $|\mathbf{x}|\to\infty$ (we denote by $\mathbf{x}$ the spatial components of $x$, at fixed time); this requirement is needed in order to disentangle the local transformations from the global ones, that are associated to the conserved charges. Infinitesimal gauge transformations, in particular, have to satisfy $\chi^{a}(\mathbf{x})\to 0$ as $|\textbf{x}|\to\infty$. If we now denote by $\Psi[\mathbf{A}]$ the wavefunction of the system (with A denoting collectively the spatial components $A_{j}(\mathbf{x})$ for a fixed time), the variation of $\Psi[\mathbf{A}]$ due to an infinitesimal gauge transformation is (integrating by part and using $P^{a\,j}(x)=-\mathrm{i}\frac{\delta}{\delta A_{j}^{a}(x)}$)
@@
Since $\chi^{a}(\mathbf{x})$ approaches zero for large x we could integrate by part in the second step of the previous equation, and the final integral is well defined. We thus see that the constraint in Eq. (16) implements gauge invariance under infinitesimal local gauge transformations. If matter is coupled to the gauge field, a further term appears in Eq. (16), related to the variation with respect to $A^{0}$ of the matter part of the Lagrangian density. The resulting constraint once again implements gauge invariance, this time under infinitesimal local gauge transformations of both gauge and matter fields.
@@
We are now almost ready to adapt the reasoning carried out in Eq. (9) to the case of four-dimensional gauge theories. The last ingredient needed is a relation between the winding number $N(\Omega)$ and four-dimensional integrals. To this purpose we introduce the (non gauge invariant) Chern–Simons current
$j_{\scriptscriptstyle{\rm CS}}^{\mu}=\frac{g^{2}}{16\pi^{2}}\epsilon^{\,\mu\alpha\beta\gamma}\mathrm{Tr}\left(F_{\alpha\beta}A_{\gamma}+\frac{2\mathrm{i}g}{3}A_{\alpha}A_{\beta}A_{\gamma}\right)\ ,$ (23)
@@
$q\equiv\partial_{\mu}j_{\scriptscriptstyle{\rm CS}}^{\mu}=\frac{g^{2}}{16\pi^{2}}\mathrm{Tr}\left(F_{\mu\nu}\tilde{F}^{\mu\nu}\right)=\frac{g^{2}}{32\pi^{2}}\epsilon^{\,\mu\nu\alpha\beta}F^{a}_{\mu\nu}F^{a}_{\alpha\beta}\ ,$ (24)
@@
$W(\mathbf{A})=\int j_{\scriptscriptstyle{\rm CS}}^{0}(t,\textbf{x})\mathrm{d}^{3}\textbf{x}\ .$ (25)
Using the four-dimensional divergence theorem and assuming the spatial components of the Chern–Simons current to be trivial at spatial infinity (e. g. due to some fixed boundary conditions) we have
@@
$\left.\langle\mathbf{A}_{f},t_{f}|\textbf{A}_{i},t_{i}\rangle\right|_{\theta}$ $=\int_{0}^{2\pi}\mathrm{d}\theta^{\prime}\frac{1}{2\pi}\sum_{\nu=-\infty}^{\infty}\mathrm{e}^{-\mathrm{i}\nu(\theta-\theta^{\prime})}\sum_{n}\Psi_{\theta^{\prime},n}(\mathbf{A}_{f})\Psi_{\theta^{\prime},n}(\textbf{A}_{i})\mathrm{e}^{-\mathrm{i}E_{n}^{(\theta^{\prime})}(t_{f}-f_{i})}=$ (28)
$=\frac{1}{2\pi}\sum_{\nu=-\infty}^{\infty}\mathrm{e}^{-\mathrm{i}\nu\theta}\int_{0}^{2\pi}\mathrm{d}\theta^{\prime}\sum_{n}\Psi_{\theta^{\prime},n}(\mathbf{A}_{f}^{\Omega_{1}^{\nu}})\Psi_{\theta^{\prime},n}(\textbf{A}_{i})\mathrm{e}^{-\mathrm{i}E_{n}^{(\theta^{\prime})}(t_{f}-f_{i})}=\sum_{\nu=-\infty}^{\infty}\mathrm{e}^{-\mathrm{i}\nu\theta}\langle\mathbf{A}_{f}^{\Omega_{1}^{\nu}},t_{f}|\textbf{A}_{i},t_{i}\rangle=$
$=\frac{1}{2\pi}\sum_{\nu=-\infty}^{\infty}\mathrm{e}^{-\mathrm{i}\nu\theta}\int_{\mathbf{A}(t_{i})=\textbf{A}_{i}}^{\textbf{A}(t_{f})=\textbf{A}_{f}^{\Omega_{1}^{\nu}}}\mathcal{D}A(t,x)\,\exp\left(\mathrm{i}\int_{t_{i}}^{t_{f}}\mathrm{d}t\int\mathrm{d}x\mathcal{L}\right)\propto\int_{\mathbf{A}(t_{i})=\textbf{A}_{i}}^{\textbf{A}(t_{f})=\textbf{A}_{f}}\mathcal{D}A(t,x)\,\exp\left(\mathrm{i}\int_{t_{i}}^{t_{f}}\mathrm{d}t\int\mathrm{d}x\left[\mathcal{L}+\theta q\right]\right)\ ,$
@@ 2 The strong CP problem
The Lagrangian density of QCD, taking into account also the matter fields and the $\theta$-term, is thus
@@
where $D_{\mu}=\partial_{\mu}-\mathrm{i}gA_{\mu}^{a}T^{a}$ is the covariant derivative in the fundamental representation of the gauge group and “f” is a flavor index. It is well known that the first two terms of this Lagrangian density (just like in Quantum Electrodynamics) are invariant under both parity (P) and charge conjugation (C), however the $\theta$-term explicitly breaks P and CP for generic values of the $\theta$ angle, and this is essentially due to the presence of the Levi-Civita totally antisymmetric tensor. If we introduce the three dimensional chromoelectric and chromomagnetic fields in analogy with classical electromagnetism we have indeed that $q\propto\mathbf{E}^{a}\cdot\mathbf{B}^{a}$, and since the chromomagnetic field is even under parity transformation we see that the $\theta$-term is odd under both P and CP. Invariance under parity and CP is clearly recovered for $\theta=0$, but also for $\theta=\pi$: since $\theta$ is defined up to $2\pi$, we have indeed $\pi\sim-\pi$. Note that at $\theta=\pi$ the discrete symmetries are however expected to be spontaneously broken, since at $\theta=\pi$ a level crossing happens, just like in Eq. (11) (see also the following discussion on $\theta$-dependence in chiral perturbation theory in Sec. 2 and Witten:1980sp ; Gaiotto:2017yup ).
@@
To discuss why this is not possible when considering the whole Standard Model we have to introduce one last ingredient: the chiral anomaly Adler:1969gk ; Bell:1969ts . If we consider a four-dimensional gauge theory coupled to a single flavor of quark, the chiral rotation
@@
This relation implies that we can reabsorb $\theta$ in the mass term at the expense of using the complex mass $m\mathrm{e}^{-\mathrm{i}\theta}$. Note that this does not change the invariance properties of the theory, since a complex mass is not invariant under P and CP transformation, but gives us the possibility of studying $\theta$-dependence using chiral perturbation theory if light quarks are present (and this is the rationale behind Eq. (31), since $m_{\pi}^{2}\propto m$). An immediate consequence of this fact is that there is no $\theta$-dependence at all if $m=0$, since in this case we can shift $\theta$ at will without affecting the mass term of the Lagrangian. When both the $\theta$-term and a complex mass are present, we can define the physical $\theta$ angle as the combination $\theta_{\scriptscriptstyle{\rm phys}}=\theta-\arg m$, which is invariant under chiral transformations and corresponds to the value of the $\theta$ angle when the fermion mass is real and positive. The previous discussion can be generalized to the case of a generic number of flavors with mass term
@@
where $M$ is a $N_{{\scriptscriptstyle{\mathrm{f}}}}\times N_{{\scriptscriptstyle{\mathrm{f}}}}$ matrix and the physical $\theta$ angle is in this case given by
$\theta_{\scriptscriptstyle{\rm phys}}=\theta-\arg\det M\ .$ (37)
The previous statement regarding the upper limit of the $\theta$ angle should thus more precisely be written as $|\theta_{\scriptscriptstyle{\rm phys}}|\lesssim 10^{-10}$, and the presence of $\theta_{\scriptscriptstyle{\rm phys}}$ shows why we can not simply fix to zero the $\theta$ angle in the Standard Model: the mass matrix originates from the electroweak sector, and since in the electroweak sector P and CP are not conserved, there is no way to fix $\theta_{\scriptscriptstyle{\rm phys}}=0$ using symmetry arguments. The fact that $|\theta_{\scriptscriptstyle{\rm phys}}|$ is so small or, equivalently, the absence of P and CP violations in the strong sector of the Standard Model, constitutes the so called strong CP problem. As noted before there is no $\theta$-dependence at all if at least one quark is massless ($\arg\det M$ is not defined if $M$ is singular). In the past, a massless up quarks has thus been suggested as a simple solution of the strong CP problem, however this is today known to be inconsistent with experimental and lattice QCD results FlavourLatticeAveragingGroupFLAG:2021npn ; Sanfilippo:2015era ; ParticleDataGroup:2024cfk (see also Alexandrou:2020bkd ).